At what temp the velocity of O2 molecules have same velocity as SO2??

Grade:12

1 Answers

Vikas TU
7 years ago
Vrms for O2 => root(3*R*T1/32)
and 
Vrms for SO2 => root(3*R*T2/64)
Equating both the velocities we get,
T1/32 = T2/64
T1/T2 = 32/64 => ½
Hence, the temp. rlation for both the gases are => 
T2 = 2T1